OpenStudy (loganator):

PLEASE HELP I"LL GIVE MEDAL Which of the following represents a direct variation? An y table shows the following coordinate pairs: ten comma seven, thirty-five comma two, and five comma fourteen. An y table shows the following coordinate pairs: ten comma seven, thirty-five comma two, and five comma fourteen. Lance bought 3 CDs and spent forty-five dollars. Jacob bought 5 CDs and spent sixty-five dollars. y equals five-eighths x I will post the table in the comment section

OpenStudy (loganator):

here is the table

OpenStudy (mertsj):

Divide y by x. If you always get the same number, it is a direct variation.

OpenStudy (loganator):

please note that when it begins to say lance bought 3 cd's that is a different option

OpenStudy (loganator):

i did that, i still can't tell

OpenStudy (loganator):

when i divide x and y values in the table they all come out messed up on the calculator like 3.1111111111 and the other ones i did didn't make much sense either,

OpenStudy (loganator):

I'm not giving medal for that

OpenStudy (mertsj):

Why don't you just write the ordered pairs instead of something comma something

OpenStudy (mertsj):

So according to your division results, it can't be the table.

OpenStudy (mertsj):

It can't be the CD's because 45/3 is not the same as 65/5

OpenStudy (loganator):

yes just by looking at the pattern of the table, it looks inverse

OpenStudy (loganator):

it's has to be one of the other 2 and C just confuses me

OpenStudy (mertsj):

So look at the equation, y = 5/8 x y/x=5/8

OpenStudy (mertsj):

Doesn't look like that will ever change.

OpenStudy (loganator):

what/ what are you talking about

OpenStudy (mertsj):

your third option

OpenStudy (mertsj):

What are you talking about?

OpenStudy (loganator):

Omg i justed asked if you knew how to do it not a long confusing debate I THINK IT'S B

OpenStudy (loganator):

Wait you said back ther it can't be cd's but how does the other one make sense?

OpenStudy (loganator):

how does y = 5/8 x make sense

OpenStudy (mertsj):

No matter what ordered pair makes that sentence true, when you divide y by x you will always get 5/8

OpenStudy (loganator):

ok so your point is? that nakes it direct variation? whatever, thanks for the help

OpenStudy (mertsj):

Remember, at the very beginning we said if it is a direct variation, when you divide y by x you will always get the same number.

OpenStudy (loganator):

kind of

OpenStudy (mertsj):

So we found out that didn't work for the table they gave us.

OpenStudy (mertsj):

We also found out that it didn't work for the CD choice

OpenStudy (loganator):

ok then process of emimination. Thx

OpenStudy (mertsj):

That brought us to the equation y = 5/8 x. And it does work for that one.

OpenStudy (mertsj):

Also you should learn that another common form of a direct variation is y = kx where k is some number and the third one works for that also.